OpenStudy (anonymous):

How would I rewrite (a^3/4)^5 using radicals?

OpenStudy (ikram002p):

humm (a^15 / 4^5) ??

OpenStudy (anonymous):

Hmmm.. I have no idea, but how did you do that? Is there a formula?

OpenStudy (ikram002p):

yess (X^a)^b=X^(a*b)

OpenStudy (ikram002p):

(X/Y)^a= X^a/Y^a

OpenStudy (anonymous):

Oh! Okee dokee. Makes sense than.

OpenStudy (ikram002p):

using both (X^b/Y)^a= X^a*b)/Y^a

OpenStudy (ikram002p):

any way radicals = root

OpenStudy (anonymous):

would I be following the same process rewriting this as a rational expononent?

OpenStudy (ikram002p):

so i thing it will goes like this(but not sure) (a^3/4)^5 =(a^15 / 4^5) =(a^15 / 4^5)^1..... but 1=2/2 =(a^15 / 4^5)^2/2 =(a^30/4^10)^1/2 sqrt (a^30/4^10)

OpenStudy (ikram002p):

yess i guss :) but there is many formulas u cud use

OpenStudy (anonymous):

Oh Okay! Great. Thanks

OpenStudy (ikram002p):

wlc :)

OpenStudy (anonymous):

Does x^2/5 sound right?

OpenStudy (ikram002p):

humm s.t x= ?

OpenStudy (anonymous):

I'm confused... Is that a formula?

OpenStudy (ikram002p):

u can convert to x^2/5 for sure , but u need to be carfull how to get X

OpenStudy (anonymous):

Ah. okay gotcha. Yay!

OpenStudy (ikram002p):

lets try this (a^3/4)^5 let it all be Y so how can u convert Y to X^2/5 ? Y^1 but 1 = 2*5/5*2 right ? so take 2/5 (Y^(2/5))^(5/2) right ?? so let (Y^(2/5))=X X^(5/2) got it ??

OpenStudy (anonymous):

Ah yes. Thank you so much for the explanation!!

OpenStudy (anonymous):

You really should teach math! You're good at explaining and showing diagrams!

OpenStudy (ikram002p):

r u getting it ?

OpenStudy (anonymous):

Yes. Much more than before

OpenStudy (ikram002p):

ohk i hope im not confusing u :o gud luck :)

OpenStudy (anonymous):

I was getting a little lost, but I really appreciate you trying to explain it! I got most it though, so thanks :)

OpenStudy (ikram002p):

np :D